solve for n if P(n,3) = 24C(n,4)
Answer by stanbon(75887)

solve for n if P(n,3) = 24C(n,4)
--------
n!/(n-3)! = 24
------------
n(n-1)(n-2) = 24 * [n(n-1)(n-2)(n-3)/24]
---------
n(n-1)(n-2) = [n(n-1)(n-2)(n-3)]
-----
n-3 = 1
n = 4
----
Checking::
P(4,3) = 24*C(4,4)
4!/1! = 24*1
---
24 = 24
---------------
cheers,
Stan H.
-----------